Summary of DELANA processings in 1998
DELANA processings in 1998
Below is a summary of the main Delana processings made during 1998 .
Early incomplete processings of the data and the numerous processings
of calibration data used to prepare full reprocessings or to test code
developments are not included.
For all processings during 1998, the
New Forward Tracking (NFT) version of the code was used.
The 98C processing was started on June 17th. Data taken prior to this
date were reprocessed and the processing was continued until the end
of the data taking.
Using the data taken during cosmic running and the B processing of the
first period of Z0 peak data, the following alignment and detector
calibrations were available from the start of the C processing.
During the course of the C processing , updates were allowed to the
code and data base where these were necessary to prepare the D
processing. The main updates were :
- VD internal alignment.
- VFT internal alignment and mapping of noisy channels.
- Corrections for T0 of TPC and first distortion coefficients for 98.
- Timing corrections for ID plus preliminary fix routines for 98.
- FCB internal calibrations.
- Global alignment of tracking chambers.
- Energy calibration of HPC from 1st Kr run, and approximate Z
- FEMC energy calibration.
- Barrel RICH calibration and short corrections.
- The same error scaling factors as used for 97E were applied.
Known problems with the C processing are:
- Jul 2 - New internal VD alignment and new global alignment.
- Jul 15 - Second set of TPC distortions based on 189 GeV data.
- Aug 3 - Correction for cable swapping in VFT readout (from run 85977).
- Aug 26 - New internal alignment of VFT and global alignment of forward region.
- Sep 6 - Shorts calibration for barrel RICH (** see below **).
- Sep 9 - Internal calibration of FCB.
- Sep 11 - Internal calibration of FCA.
- The update for the shorts of the Barrel RICH made on Sep 6th was
corrupted. Barrel RICH data is unreliable in the C processing for
runs after this date.
- A cable swap affecting sector 4 pad row 8 of the TPC was corrected
in the hardware on 23rd June. This pad row is essentially 'off' for
data prior to this date, and the distortion coefficients for this
row are poorly determined.
- As noted above , VFT data prior to Aug 3rd are affected by cable swaps.
The 98D processing is the first fully consistent processing of this
Note that this processing is the first to use the new pattern recognition
in the Hadron Calorimeter based of the cathode read-out.
In addition, futher refinements to the tracking algorithms were introduced.
The data were processed in three batches
Compared with the situation at the END of the 98C processing, the additional updates were:
- 98Z_D proc 4 Nov 1998 ---> 17 Nov 1998 (1st period of Z0 data)
- 98H_D proc 13 Nov 1998 ---> 5 Dec 1998 (189 GeV data)
- 98Z_D proc 4 Dec 1998 ---> 6 Dec 1998 (2nd period of Z0 data)
- Sofware correction for the TPC cable swap in the data before June 23rd and
update of the TPC distortion coefficients.
- Full IDFIX for the 98 data.
- Correction for cabling problem in VFT, and for a shift in the min-strip
readout affecting the early Z0 data.
- Energy calibration and alignment of HPC for all 98 data, and correction
of a problem affecting the energy of showers in the outer Z crack.
- Correction of the RICH data base problem and calibration of shorts until
the end of data taking.
- FEMC energy calibration extended to the end of data taking.
- The Background 1 level is recorded in the pilot record as an additional
word in the LEP blocklet (stored as NINT(1000.*BKG1)).
Comments/questions to John.Wickens@cern.ch
Last update Dec 17 1998